# A Practitioner’s Guide to Kolmogorov–Arnold Networks

### v1

@@ 12 Current Gaps and Path Forward
The rapid proliferation of KANs, marked by an explosion of publications and open-source implementations (Tables 2 and 3), signals a field rich with opportunity yet characterized by fragmentation. While empirical studies have demonstrated KANs’ potential across regression, operator learning, and PDE solving, their comparative evaluations remain inconsistent. The field now stands at a crossroads: rather than engaging in superficial “KAN vs. MLP” contests, the next phase must focus on building a rigorous, principled science of KAN design, training, and theory.
The Limits of Current Comparative Frameworks
Most current KAN literature follows one of two dominant paradigms, both of which—though valuable—have clear limitations:
(1) Superficial “KAN vs. MLP” Benchmarks.
Many studies pit a single KAN variant against a baseline MLP with matched parameter counts. However, this comparison is often ill-defined: KANs and MLPs differ in the functional roles of parameters, the implicit regularization introduced by basis structures, and the stability constraints imposed by normalization or grid design. Consequently, results vary widely, from strong KAN outperformance to near equivalence [28, 29, 30]. Without a shared theoretical grounding, such results offer limited insight into why or when a KAN excels.
(2) Problem-Dependent “Basis-vs.-Basis” Races.
Another thread compares various bases—splines, Chebyshev polynomials, Gaussians—across selected problems [31]. While these studies show that performance is problem-dependent, they risk obscuring deeper principles. Each basis family spans a vast design space of normalization schemes, grid choices, and regularization methods. A single benchmark cannot meaningfully represent such diversity.
@@
In sum, existing comparisons—whether architecture-based or basis-based—provide no predictive framework for design. A KAN is not a monolithic model; it is a framework whose performance depends on principled design choices rooted in numerical analysis, approximation theory, and optimization geometry.
Case Study: The Many Faces of a Chebyshev KAN
The “Chebyshev KAN” exemplifies both the pitfalls of naïve design and the power of principled refinement. A straightforward Chebyshev expansion often exhibits unstable training and inconsistent results. However, as multiple studies demonstrate, careful modifications dramatically improve stability and expressivity:
@@
  • Domain stabilization: Applying tanh normalization at the input and between layers constrains arguments to [1,1][-1,1], where Chebyshev polynomials are numerically stable (Section 6.2).
  • Architectural hybrids: For inverse and PDE problems, replacing the final Chebyshev expansion with a linear head reduces overfitting and improves conditioning [57].
@@
  • Grid-informed adaptivity: Introducing a small, learnable tanh grid before expansion, as in [44], provides local adaptivity and stabilizes training.
  • Principled optimization: Hybrid Adam–LBFGS schedules and modest weight decay help navigate sharp spectral loss landscapes [56, 55].
@@
Once these refinements are incorporated, Chebyshev-based KANs become powerful and interpretable approximators. Similar trends hold for other bases: Gaussian KANs improve via adaptive centers [46]; B-spline KANs through free-knot adaptation [77, 38]; and finite-basis KANs through domain decomposition [86].
A Multi-Pillar Framework for the Next Generation of KAN Research
We propose a unified, multi-stage research agenda structured around six complementary pillars. Together, they chart a roadmap from empirical engineering to a comprehensive theoretical science of KANs.
Pillar 1: Building a Robust Component Library.
The immediate task is to consolidate best practices into a library of well-engineered, stable KAN layers. Each major basis family (Chebyshev, spline, Gaussian, Fourier, wavelet) requires a canonical, reproducible implementation with optimized normalization, initialization, and regularization. This engineering groundwork establishes reliable building blocks—such as StableChebyKANLayer or AdaptiveSplineKANLayer—for systematic study and fair benchmarking.
Pillar 2: Developing a Science of Basis Selection.
Basis choice is the defining inductive bias of a KAN, yet current guidelines are purely heuristic. A formal basis selection theory should link the mathematical properties of the target problem to the optimal basis type. Key open questions include:
  • How do spectral properties of a basis (global vs. local support) align with the Fourier spectrum of the data or PDE operator?
  • Can we prove that Chebyshev bases improve conditioning of the NTK for elliptic PDEs?
  • How can we quantify the trade-off between approximation power and implicit regularization (e.g., smooth B-splines vs. oscillatory polynomials)?
This theoretical bridge would replace trial-and-error tuning with principled, problem-dependent design.
Pillar 3: Toward a Comprehensive Optimization Theory Beyond the NTK Regime.
Existing convergence guarantees for KANs mirror those of MLPs under the Neural Tangent Kernel (NTK) framework [138, 56]. While powerful, NTK theory relies on assumptions that do not hold in practice—namely infinite width and “lazy training” near initialization. Real KANs are finite and undergo significant parameter movement, particularly in their learnable basis parameters (e.g., knot locations, RBF centers). A new optimization theory must therefore:
  1. 1.
    Describe the geometry of the loss landscape for finite-width KANs.
  2. 2.
    Characterize conditions that prevent convergence to poor local minima.
  3. 3.
    Analyze how adaptive basis parameters reshape this landscape during feature learning.
Such a theory would bridge the gap between NTK-based asymptotics and the empirical reality of nonlinear, feature-adaptive training.
Pillar 4: Understanding Generalization and Regularization.
KANs blur the boundary between architecture and regularization: the choice of basis implicitly constrains the function class. A low-degree polynomial basis naturally suppresses high-frequency noise, while a fine B-spline grid enables high-resolution fitting. Future work must quantify how this implicit bias interacts with explicit regularizers (e.g., L1, L2, entropy). Developing generalization bounds that depend jointly on the number of parameters and basis complexity (e.g., polynomial degree or grid resolution) would provide a rigorous explanation for KANs’ empirically strong generalization relative to MLPs of comparable size.
Pillar 5: A Formal Theory of Composition and Hierarchy.
While the Kolmogorov–Arnold theorem describes a two-layer representation, modern KANs are deep. We lack a theory describing how functional properties evolve under composition. Open questions include:
  • What is the functional meaning of “depth” in a KAN?
  • How do layers with heterogeneous bases (e.g., Fourier followed by spline) interact?
  • Under what conditions does composition enhance expressivity without destabilizing optimization?
A compositional function theory for deep KANs would formalize how information is processed in function space, providing principles for designing heterogeneous and hierarchical architectures.
Pillar 6: A Theory of Interpretability and Identifiability.
KANs are prized for interpretability: their univariate activations often resemble symbolic expressions like $\sin(x)$ or $x^{2}$. Yet this interpretability is empirical, not theoretical. Two different sets of learned functions can yield the same global mapping, raising issues of non-uniqueness and identifiability. A rigorous interpretability theory should:
  • Define when a learned function is a true, unique representation of the underlying law.
  • Provide conditions for stability—ensuring that small data perturbations do not radically alter learned symbolic forms.
  • Establish identifiability criteria guaranteeing that interpretable decompositions correspond to genuine structures.
This would ground symbolic interpretability in mathematical guarantees rather than anecdotal observation.
@@ A Multi-Pillar Framework for the Next Generation of KAN Research
The path forward for KANs is not a race to find a single dominant architecture. It is a structured progression: from building reliable components, to uncovering theoretical principles, to integrating them into interpretable, adaptive, and compositional systems. By developing robust engineering standards, principled basis theory, finite-width optimization frameworks, and formal interpretability guarantees, the KAN community can move beyond benchmarking toward a predictive, unified science of functional learning.
@@ 13 Conclusion
Kolmogorov–Arnold Networks have emerged as a compelling architectural paradigm, shifting the locus of learnable parameters from linear weights to univariate basis functions on the network’s edges. This review has synthesized the rapidly expanding KAN landscape, with a focus on scientific machine learning applications such as function approximation, regression, and PDE solving. Our analysis shows that KANs, while not a universal replacement for MLPs, offer a powerful and highly configurable alternative. The central theme of our findings is that the question is not whether KANs are better than MLPs, but rather which KAN is best suited for a given problem.
Across the literature, a consistent pattern emerges: well-chosen KAN variants frequently outperform vanilla MLP and PINN baselines in accuracy and convergence speed, though often at the cost of higher computational overhead per epoch. The key to unlocking this performance lies in navigating the vast design space that KANs open up. We have systematically organized this space, covering the diverse families of basis functions—from the locality of B-splines and Gaussians to the global spectral properties of Chebyshev and Fourier series—and detailing how choices in normalization, grid structure, and initialization critically impact stability and expressivity.
Furthermore, we have structured the ecosystem of techniques for enhancing KAN performance. For accuracy, strategies such as physics-informed losses, adaptive sampling, domain decomposition, and hybrid architectures have proven effective. For efficiency, innovations in GPU-aware engineering, matrix optimizations, and the use of computationally cheaper bases like ReLU-powers are mitigating KANs’ primary limitation. These advancements, coupled with principled regularization and a growing body of theoretical results on approximation rates and convergence, signal a maturing field.
The practical “Choose–Your–KAN” guide presented in Section 11 operationalizes these insights, providing a decision-making framework that maps problem characteristics to specific architectural and training choices. As we have argued, the true strength of the KAN framework lies not in a single canonical implementation, but in its modularity. By providing the tools to select, combine, and adapt basis functions, KANs empower researchers to inject valuable inductive biases tailored to the problem at hand, bridging the gap between versatile deep learning models and classical approximation theory. The journey ahead will involve deeper, basis-specific studies and standardized benchmarking, but it is clear that KANs have already secured their place as a vital and promising branch of neural network design.

### v2

@@ 12 Current Gaps and Path Forward
Despite rapid growth in publications and open-source implementations (Tables 2, LABEL:githubs), the KAN ecosystem remains fragmented. Most studies show strong empirical potential in regression, operator learning, and PDE solving, yet comparisons are inconsistent and rarely grounded in theory. The field now requires principled foundations rather than isolated “KAN vs. MLP” or “basis-vs.-basis” experiments.
Limits of Current Comparative Frameworks
(1) KAN vs. MLP benchmarks.
These comparisons often match parameter counts but ignore essential differences: functional roles of parameters, basis-induced inductive bias, and normalization/grid constraints. Outcomes range from clear KAN advantages to near parity [36, 37, 38], offering little predictive insight into when KANs should win.
(2) Basis-vs.-basis races.
Comparisons of splines, Chebyshev, Gaussians, etc. [39] highlight problem dependence but mask deeper structure. Each basis involves choices of normalization, grids, and regularization—too many degrees of freedom for a single benchmark to be conclusive.
@@
Overall, KANs are not a single architecture but a framework whose performance depends on principled design choices rooted in approximation theory, numerical conditioning, and optimization geometry.
Case Study: The Many Faces of a Chebyshev KAN:
A plain Chebyshev KAN can be unstable, but targeted refinements substantially improve behavior:
@@
  • Domain stabilization: tanh normalization confines inputs to [1,1][-1,1] for numerically stable Chebyshev evaluation (Section 6.2).
  • Architectural hybrids: Linear heads reduce overfitting in inverse/PDE tasks [65].
@@
  • Grid-informed adaptivity: Learnable tanh-based grids add local adaptivity [52].
  • Principled optimization: Hybrid Adam–L-BFGS schedules improve spectral conditioning [64, 63].
  • Adaptive polynomial degree: Making the polynomial degree itself trainable, as in Adaptive PolyKAN [280].
@@
Similar refinements benefit other bases: adaptive centers for Gaussian KANs [54], free-knot updates for spline KANs [113, 46], and domain decomposition for finite-basis KANs [122].
A Multi-Pillar Framework for Future KAN Research
Pillar 1: Robust component library.
Canonical, stable implementations for Chebyshev, spline, Gaussian/RBF, Fourier, and wavelet bases with standardized normalization and initialization.
Pillar 2: Basis selection theory.
Principles linking problem structure to basis choice, including spectral alignment, NTK conditioning, and approximation–regularization trade-offs.
Pillar 3: Finite-width optimization theory.
Beyond NTK predictions [165, 64]; characterize finite-width loss geometry, adaptive basis movement, and optimization dynamics. Include Extreme Learning Machine–style regimes [278] as an alternative training strategy for KANs.
Pillar 4: Generalization and regularization.
Joint treatment of parameter count, basis complexity, smoothness, and explicit $\ell_{1}/\ell_{2}$ penalties for improved generalization bounds.
Pillar 5: Composition and hierarchy.
Depth in KANs currently has no formal interpretation. Open problems include understanding the functional role of layer composition, the interactions among heterogeneous bases, and the trade-off between expressivity and stability under deep compositions (see also the theoretical results of [53, 273] and the empirical observations of [267] [^51] https://github.com/geoelements-dev/mlp-kan).
Pillar 6: Interpretability and identifiability.
Uniqueness of learned univariate functions, robustness, and symbolic identifiability.
Pillar 7: RKHS-based formulation.
Develop RKHS-guided KAN theory using recent formulations [273] and earlier functional-analytic perspectives on KST [100].
Pillar 8: New basis functions.
Introduce and test new basis families in KANs, including Newton basis functions [277, 279], low-order Wendland functions, and Matérn/Sobolev kernels [275].
@@ A Multi-Pillar Framework for Future KAN Research
A unified framework combining stable components, principled basis selection, finite-width theory, RKHS insights, and exploration of new bases can guide future KAN development.
@@ 13 Conclusion
Kolmogorov–Arnold Networks shift the locus of learning from linear weights to univariate basis functions placed on edges, offering a flexible alternative to conventional MLPs. This review has consolidated the rapidly expanding KAN literature, with emphasis on applications in function approximation, regression, and PDE solving. Our central observation is that the key question is not whether KANs outperform MLPs, but rather which KAN is appropriate for a given task.
A clear pattern emerges across studies: well-designed KAN variants often exceed vanilla MLP and PINN baselines in accuracy and convergence, albeit with higher per-epoch computational cost. Unlocking this performance hinges on navigating KANs’ large design space. We organized this space through the lens of basis families—local B-splines and Gaussians, global Chebyshev and Fourier series—and highlighted how normalization, grid layout, and initialization shape stability and expressivity.
We also outlined the ecosystem of techniques for enhancing KAN performance. Accuracy gains arise from physics-informed losses, adaptive sampling, domain decomposition, and hybrid sequence/attention modules. Efficiency gains are enabled by GPU-aware engineering, matrix-level optimizations, and lighter bases such as ReLU-powers. These developments, combined with emerging theory on approximation and convergence, signal a transition from exploratory experimentation to principled model design.
The practical “Choose–Your–KAN” guide in Section 11 distills these insights into a decision framework linking problem characteristics to architectural and training choices. Ultimately, the strength of the KAN framework lies in its modularity: basis families, normalization schemes, and optimization strategies can be selected and composed to encode problem-specific inductive biases. Continued progress will require deeper basis-specific analysis and standardized benchmarks, but KANs have clearly established themselves as a versatile and promising direction in neural network architecture.